Wind resistance or atmospheric drag tends to slow down moving objects. Atmospheric drag varies jointly as an object's area "A" and velocity "v". If a car traveling at a speed of 50 mph with a surface area of 35Ftsquared experiences a drag of 192.5N (Newton’s), how fast must a car with 49ft squared surface area travel in order to experience a drag force of 247.94N?
A. 46 mph
B. 43 mph
C. 51 mph
D. 48 mph
B. 43 mph
C. 51 mph
D. 48 mph

Answer by venugopalramana(3286)

LET DRAG FORCE BE=F ,AREA=A,AND VELOCITY=V
F VARIES JOINTLY AS A AND V.THAT MEANS IT VARIES AS A*V
HENCE F=K*A*V…WHERE K IS A CONSTANT.
WE GOT FOR CASE 1….F=192.5=K*35*50..HENCE K=192.5/(50*35)=0.11
HENCE FOR CASE 2 ….F=247.94=0.11*49*V….V=247.94/(0.11*49)=46……A IS THE ANSWER